# Gibbs-preserving operations requiring infinite amount of quantum coherence

### v2

@@ Appendix D State transformations requiring cost-diverging Gibbs preserving Operations
${\cal R}(...)=\mathrm{Tr}[|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|...]|\eta_{+}\rangle\langle\eta_{+}|_{S}+\mathrm{Tr}[(1-|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|)...]|\eta_{-}\rangle\langle\eta_{-}|_{S},$ (84)
which perfectly recovers $|\eta_{\pm}\rangle$ because of (83). We can also directly check that ${\cal C}(\Lambda,\mathbb{P})>0$ as follows. Note that the operator $\Lambda^{\dagger}(H^{\prime})$ satisfies
$\langle\eta_{+}|_{S}\Lambda^{\dagger}(H^{\prime})|\eta_{-}\rangle_{S}$ $=\sum_{k}\langle\eta_{+}|_{S}M^{\dagger}_{k}H^{\prime}M_{k}|\eta_{-}\rangle_{S}$
@@
${\cal C}(\Lambda,\mathbb{P})^{2}$ $=|\langle\eta_{+}|H|\eta_{-}\rangle_{S}|^{2}$
@@
where the last inequality is because of the assumption $\tau_{S,i}<\tau_{S^{\prime},i^{\prime}}<\tau_{S,j}$, which ensures $E_{S,i}-E_{S,j}\neq 0$ and $10<r<$ 1. Applying Theorem 2 to $\Lambda$ and $\mathbb{P}$ results in ${\cal F}_{c}^{\epsilon=0}(\Lambda)=\infty$, concluding the proof. ∎
Appendix E Upper bound for the coherence cost of Eq. (2)
satisfies the upper bound
${\cal F}_{c}(\Lambda)\leq{\cal F}(\eta)+{\cal F}(\sigma).$ (88)
Proof.
$\Lambda_{1}(\kappa_{S}):=\langle 1|\kappa_{S}|1\rangle|{1}\rangle\!\langle{1}|_{E}+\langle 0|\kappa_{S}|0\rangle|{0}\rangle\!\langle{0}|_{E}$ (89)
@@
$\Lambda_{2}(\kappa_{E}):=\langle 1|\kappa_{E}|1\rangle\eta+\langle 0|\kappa_{E}|0\rangle\sigma$ (90)
for arbitrary states $\kappa_{S}$ in $S$ and $\kappa_{E}$ in $E$.
$\Lambda_{1}(\kappa_{S})$ $=\mathrm{Tr}_{S}[U\kappa_{S}\otimes|0\rangle\langle 0|_{E}U^{\dagger}]$ (91)
$\Lambda_{2}(\kappa_{E})$ $=\mathrm{Tr}_{E\tilde{S}}[V\kappa_{E}\otimes\eta\otimes\sigma V^{\dagger}]$ (92)
where
$U$ $:=|00\rangle\langle 00|_{SE}+|11\rangle\langle 10|_{SE}+|01\rangle\langle 01|_{SE}+|10\rangle\langle 11|_{SE},$ (93)
$V$ $:=|0\rangle\langle 0|_{E}\otimes\mathds{1}_{S\tilde{S}}+|1\rangle\langle 1|_{E}\otimes U_{\mathrm{SWAP}}$ (94)
${\cal F}_{c}(\Lambda_{1})$ $\leq{\cal F}(|{0}\rangle\!\langle{0}|)=0$ (95)
${\cal F}_{c}(\Lambda_{2})$ $\leq{\cal F}(\eta\otimes\sigma)={\cal F}(\eta)+{\cal F}(\sigma).$ (96)
Since ${\cal F}_{c}(\Lambda)\leq{\cal F}_{c}(\Lambda_{1})+{\cal F}_{c}(\Lambda_{2})$, we obtain (88). ∎
Appendix F Upper bounds (Proofs of Theorems 5 and 6)
@@
$\sqrt{{\cal F}_{c}(\Lambda_{\epsilon})}\leq\sqrt{{\cal F}_{c}({\cal V}_{\epsilon})}\leq\frac{\Delta(H_{\rm tot}-V^{\dagger}H_{\rm tot}V)}{2\epsilon}+\sqrt{2}\Delta(H_{\rm tot}).$ (99)
@@
$D_{F}(\Lambda_{\epsilon},\Lambda)$ $=D_{F}(\mathrm{Tr}_{E^{\prime}}\circ{\cal V}_{\epsilon}\circ{\cal P}_{|\eta\rangle},\mathrm{Tr}_{E^{\prime}}\circ{\cal V}\circ{\cal P}_{|\eta\rangle})$ (100)
@@
where the second line comes from the definition of the channel purified distance (25), the third line is because of the data-processing inequality of the purified distance, and the fourth line follows from the assumption of ${\cal V}_{\epsilon}$. ∎
@@
$\Lambda(\rho)=\langle{+}|{\rho}|{+}\rangle|{0}\rangle\!\langle{0}|+\langle{-}|{\rho}|{-}\rangle|{1}\rangle\!\langle{1}|.$ (102)
This is evidently a Gibbs-preserving Operation, noting that $H_{S^{\prime}}=0$. Take the state pair $\mathbb{P}=\{|{+}\rangle\!\langle{+}|,|{-}\rangle\!\langle{-}|\}$. Since both states are reversible under $\Lambda$, we have $\delta(\Lambda,\mathbb{P})=0$, and ${\cal C}(\Lambda,\mathbb{P})>0$ can be checked by direct computation. This ensures that Theorem S.2 can be applied, and the lower bound in (101) then immediately follows noting that $\Delta(H_{S})=\tilde{a}\leq a$ and $H_{S^{\prime}}=0$.
@@
$\Lambda(\rho)=\mathrm{Tr}_{S}(V\rho\otimes|{0}\rangle\!\langle{0}|_{S^{\prime}}V^{\dagger})$ (103)
@@
$H_{\mathrm{tot}}-V^{\dagger}H_{\mathrm{tot}}V=H_{S}\otimes\mathds{1}-V^{\dagger}(H_{S}\otimes\mathds{1})V=(H_{S}-U_{H}H_{S}U_{H})\otimes\mathds{1}.$ (104)
@@
$\Delta(H_{\mathrm{tot}}-V^{\dagger}H_{\mathrm{tot}}V)=\Delta(H_{S}-U_{H}H_{S}U_{H})=\sqrt{2}\tilde{a}.$ (105)
@@
${\cal C}(\Lambda,\mathbb{P})=\left|\langle{+}|{H_{S}-U_{H}H_{S}U_{H}}|{-}\rangle\right|=\frac{\tilde{a}}{2}.$ (106)
Combining (105) and (106) gives
$\Delta(H_{\mathrm{tot}}-V^{\dagger}H_{\mathrm{tot}}V)=2\sqrt{2}\,{\cal C}(\Lambda,\mathbb{P}),$ (107)
from which the upper bound in (101) follows by using Theorem S.7 and noting $\Delta(H_{S})=\tilde{a}=a/\sqrt{2}$ and $\Delta(H_{S^{\prime}})=0$. ∎

### v3

@@ VI State transformations requiring cost-diverging Gibbs preserving Operations (Proof of Theorem 3)
${\cal R}(...)=\mathrm{Tr}[|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|...]|\eta_{+}\rangle\langle\eta_{+}|_{S}+\mathrm{Tr}[(1-|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|)...]|\eta_{-}\rangle\langle\eta_{-}|_{S},$ (S.129)
which perfectly recovers $|\eta_{\pm}\rangle$ because of (S.128). We can also directly check that ${\cal C}(\Lambda,\mathbb{P})>0$ as follows. Note that the operator $\Lambda^{\dagger}(H_{S^{\prime}})$ satisfies
$\langle\eta_{+}|_{S}\Lambda^{\dagger}(H_{S^{\prime}})|\eta_{-}\rangle_{S}$ $=\sum_{k}\langle\eta_{+}|_{S}M^{\dagger}_{k}H_{S^{\prime}}M_{k}|\eta_{-}\rangle_{S}$
@@
${\cal C}(\Lambda,\mathbb{P})^{2}$ $=|\langle\eta_{+}|H_{S}|\eta_{-}\rangle_{S}|^{2}$
@@
where the last inequality is because of the assumption $\tau_{S,i}<\tau_{S^{\prime},i^{\prime}}<\tau_{S,j}$, which ensures $E_{S,i}-E_{S,j}\neq 0$ and $10<r<$ 1. Applying Theorem 1 to $\Lambda$ and $\mathbb{P}$ results in ${\cal F}_{c}^{\epsilon=0}(\Lambda)=\infty$.
Next, we show (iii), i.e. $|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}}$ satisfies (S.111). It suffices to show that for an arbitrary state $\sigma^{\prime}_{S^{\prime},\epsilon}$ such that $D_{F}(|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}},\sigma^{\prime}_{S^{\prime},\epsilon})\leq\epsilon$, the following inequality holds:
$\sqrt{{\cal F}^{\epsilon=0}(|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow\sigma^{\prime}_{S^{\prime},\epsilon})}\geq\frac{\sqrt{4\kappa(\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}}{\sqrt{\epsilon}}-\Delta(H_{S})-3\Delta(H_{S^{\prime}}).$ (S.132)
Derivation of (S.111) from (S.132) is as follows: First, due to $(a-b)^{2}\geq a^{2}/2-b^{2}$, we obtain
${\cal F}^{\epsilon=0}(|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow\sigma^{\prime}_{S^{\prime},\epsilon})\geq\frac{2\kappa(\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}{\epsilon}-(\Delta(H_{S})+3\Delta(H_{S^{\prime}}))^{2}.$ (S.133)
Second, if a channel $\tilde{\Lambda}$ satisfies $\tilde{\Lambda}(\rho^{\prime})=\sigma^{\prime}_{S^{\prime},\epsilon}$ for a state $\rho^{\prime}$ satisfying $D_{F}(\rho^{\prime},|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S})\leq\epsilon$, the channel $\tilde{\Lambda}$ also satisfies $D_{F}(\tilde{\Lambda}(|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}),|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}})\leq 2\epsilon$. Therefore, ${\cal F}^{\epsilon}(|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}})\geq\min_{\sigma^{\prime\prime}_{S^{\prime}}:D_{F}(\sigma^{\prime\prime}_{S^{\prime}},|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}})\leq 2\epsilon}{\cal F}^{0}_{c}(|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow\sigma^{\prime\prime}_{S^{\prime}})$. Therefore, (S.111) follows (S.132).
Let us show (S.132). Let $\Lambda$ be an arbitrary Gibbs-preserving operation satisfying
$D_{F}(\Lambda(|\eta_{+}\rangle\langle\eta_{+}|_{S}),|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}})\leq\epsilon.$ (S.134)
[…]
$\sqrt{{\cal F}_{c}(\Lambda)}\geq\frac{\sqrt{4\kappa(\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}}{\sqrt{\epsilon}}-\Delta(H_{S})-3\Delta(H_{S^{\prime}}).$ (S.135)
To show this, we follow a similar strategy of proof of (ii). Let $\{M_{k}\}_{k}$ be a Kraus representation of $\Lambda$. We write each Kraus operator with respect to energy eigenbases $\{|l\rangle_{S}\}_{l=1}^{d_{S}}$ of $S$ and $\{|m\rangle_{S^{\prime}}\}_{m=1}^{d_{S^{\prime}}}$ of $S^{\prime}$ as
$+\sum_{m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{-}|_{S}+\sum_{l\neq i,j}e_{k,l}|i^{\prime}\rangle_{S^{\prime}}\langle l|_{S}+\sum_{m\neq i^{\prime}}\sum_{l\neq i,j}f_{k,m,l}|m\rangle_{S^{\prime}}\langle l|_{S},$ (S.136)
where $|\eta_{-}\rangle:=\sqrt{1-r}|i\rangle_{S}-\sqrt{r}|j\rangle_{S}$.
The assumption (S.134) imposes $\langle i^{\prime}|_{S^{\prime}}\Lambda(|\eta_{+}\rangle\langle\eta_{+}|_{S})|i^{\prime}\rangle_{S^{\prime}}\geq 1-\epsilon^{2}$, and thus we obtain
$\sum_{k}|a_{k}|^{2}\geq 1-\epsilon^{2}.$ (S.137)
Furthermore, $\sum_{k}M^{\dagger}_{k}M_{k}=\mathds{1}_{S}$ implies
11 11 $=\sum_{k}\langle\eta_{+}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=\sum_{k}|a_{k}|^{2}+\sum_{k}\sum_{m\neq i^{\prime}}|b_{k,m}|^{2},$ (S.138)
00 0 $=\sum_{k}\langle\eta_{-}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=\sum_{k}c^{*}_{k}a_{k}+\sum_{k}\sum_{m\neq i^{\prime}}d^{*}_{k,m}b_{k,m},$ (S.139)
00 0 $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=\sum_{k}e^{*}_{k,l}a_{k},\enskip\forall l\mbox{ s.t. }l\neq i,j,$ (S.140)
00 0 $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{-}\rangle=\sum_{k}e^{*}_{k,l}c_{k},\enskip\forall l\mbox{ s.t. }l\neq i,j.$ (S.141)
From these relations, we obtain
$\sum_{k}\sum_{m\neq i^{\prime}}|b_{k,m}|^{2}$ $=1-\sum_{k}|a_{k}|^{2}\leq\epsilon^{2},$ (S.142)
$|\sum_{k}c^{*}_{k}a_{k}|$ $=|\sum_{k}\sum_{m\neq i^{\prime}}d^{*}_{k,m}b_{k,m}|$
$\leq\sqrt{\sum_{k}\sum_{m\neq i^{\prime}}|d_{k,m}|^{2}}\sqrt{\sum_{k}\sum_{m\neq i^{\prime}}|b_{k,m}|^{2}},$
$\leq\epsilon.$ (S.143)
Therefore, we obtain
$=$ $\sum_{k}|a_{k}|^{2}(r\tau_{S,i}+(1-r)\tau_{S,j})+\sum_{k}|c_{k}|^{2}((1-r)\tau_{S,i}+r\tau_{S,j})$
$+\sum_{k}\sum_{l\neq i,j}|e_{k,l}|^{2}\tau_{S,l}+2\mathrm{Re}\left(\sum_{k}c^{*}_{k}a_{k}\right)\sqrt{r(1-r)}(\tau_{S,i}-\tau_{S,j})$
$\geq$ $(1-\epsilon^{2})\tau_{S^{\prime},i^{\prime}}+\sum_{k}|c_{k}|^{2}((1-r)\tau_{S,i}+r\tau_{S,j})-2\epsilon\sqrt{r(1-r)}(\tau_{S,j}-\tau_{S,i}).$ (S.144)
Because of the definition of $r$, (S.144) is converted as follows:
$\sum_{k}|c_{k}|^{2}$ $\leq\frac{\epsilon^{2}\tau_{S^{\prime},i^{\prime}}+2\epsilon\sqrt{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})}}{\tau_{S,i}+\tau_{S,j}-\tau_{S^{\prime},i^{\prime}}}$
$=\epsilon f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j}).$ (S.145)
where
$f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j}):=\frac{\epsilon\tau_{S^{\prime},i^{\prime}}+2\sqrt{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})}}{\tau_{S,i}+\tau_{S,j}-\tau_{S^{\prime},i^{\prime}}}.$ (S.146)
Now, let us show that $\Lambda$ satisfies (S.135). We define =$\mathbb{P}=\{|{\eta_{+}}\rangle\!\langle{\eta_{+}}|,|{\eta_{-}}\rangle\!\langle{\eta_{-}}|\}$ and a recovery channel
${\cal R}(...)=\mathrm{Tr}[|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|...]|\eta_{+}\rangle\langle\eta_{+}|_{S}+\mathrm{Tr}[(1-|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|)...]|\eta_{-}\rangle\langle\eta_{-}|_{S}.$ (S.147)
Then, we obtain
$D_{F}({\cal R}\circ\Lambda(\eta_{+}),\eta_{+})^{2}$ $=1-\langle i^{\prime}|_{S^{\prime}}\Lambda(\eta_{+})|i^{\prime}\rangle_{S^{\prime}}$
$=1-\sum_{k}|a_{k}|^{2}$
$\leq\epsilon^{2},$ (S.148)
$D_{F}({\cal R}\circ\Lambda(\eta_{-}),\eta_{-})^{2}$ $=1-\mathrm{Tr}[(1-|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}})\Lambda(\eta_{-})]$
$=\sum_{k}|c_{k}|^{2}$
$\leq\epsilon f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})$ (S.149)
@@
$\delta(\Lambda,\mathbb{P})$ $\leq\sqrt{\frac{D_{F}({\cal R}\circ\Lambda(\eta_{+}),\eta_{+})^{2}+D_{F}({\cal R}\circ\Lambda(\eta_{-}),\eta_{-})^{2}}{2}}$
$\leq\sqrt{\frac{\epsilon^{2}+\epsilon f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}{2}}.$ (S.150)
We can also directly estimate ${\cal C}(\Lambda,\mathbb{P})$ as follows. Note that the operator $\Lambda^{\dagger}(H_{S^{\prime}})$ satisfies
$|\langle\eta_{+}|_{S}\Lambda^{\dagger}(H_{S^{\prime}})|\eta_{-}\rangle_{S}|$ $=|\sum_{k}\langle\eta_{+}|_{S}M^{\dagger}_{k}H_{S^{\prime}}M_{k}|\eta_{-}\rangle_{S}|$
$=\left|\sum_{k}\left(a^{*}_{k}\langle i^{\prime}|_{S^{\prime}}+\sum_{m:m\neq i^{\prime}}b^{*}_{k,m}\langle m|_{S^{\prime}}\right)H_{S^{\prime}}\left(c_{k}|i^{\prime}\rangle_{S^{\prime}}+\sum_{m:m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\right)\right|$
$=\left|E_{S^{\prime},i^{\prime}}\sum_{k}a^{*}_{k}c_{k}+\sum_{k}\sum_{m:m\neq i^{\prime}}b^{*}_{k,m}d_{k,m}E_{S^{\prime},m}\right|$
$=\left|\sum_{k}\sum_{m:m\neq i^{\prime}}b^{*}_{k,m}d_{k,m}(E_{S^{\prime},m}-E_{S^{\prime},i^{\prime}})\right|$
$\leq\sqrt{\sum_{k}\sum_{m:m\neq i^{\prime}}|b_{k,m}|^{2}}\sqrt{\sum_{k}\sum_{m:m\neq i^{\prime}}|d_{k,m}|^{2}(E_{S^{\prime},m}-E_{S^{\prime},i^{\prime}})^{2}}$
$\leq\epsilon\Delta(H_{S^{\prime}}).$ (S.151)
Therefore,
${\cal C}(\Lambda,\mathbb{P})$ $=|\langle\eta_{+}|(H_{S}-\Lambda^{\dagger}(H_{S^{\prime}}))|\eta_{-}\rangle_{S}|$
$\geq|\langle\eta_{+}|(E_{S,i}|i\rangle\langle i|_{S}+E_{S,j}|j\rangle\langle j|_{S})|\eta_{-}\rangle|-\epsilon\Delta(H_{S^{\prime}})$
$=\sqrt{r(1-r)}|E_{S,i}-E_{S,j}|-\epsilon\Delta(H_{S^{\prime}}).$ (S.152)
Applying Theorem S.1 to $\Lambda$ and $\mathbb{P}$, we obtain
$\sqrt{{\cal F}_{c}(\Lambda)}$ $\geq\frac{\sqrt{r(1-r)|E_{S,i}-E_{S,j}|}-\epsilon\Delta(H_{S^{\prime}})}{\sqrt{\frac{\epsilon^{2}+\epsilon f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}{2}}}-\Delta(H_{S})-\Delta(H_{S^{\prime}})$
$\geq\frac{\sqrt{r(1-r)|E_{S,i}-E_{S,j}|}}{\sqrt{\frac{\epsilon^{2}+\epsilon f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}{2}}}-\sqrt{2}\Delta(H_{S^{\prime}})-\Delta(H_{S})-\Delta(H_{S^{\prime}})$
$\geq\frac{\sqrt{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})|E_{S,i}-E_{S,j}|}}{(\tau_{S,j}-\tau_{S,i})\sqrt{\epsilon}\sqrt{\frac{\epsilon+f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}{2}}}-\Delta(H_{S})-3\Delta(H_{S^{\prime}})$
$\geq\frac{\sqrt{4\kappa(\epsilon,\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}}{\sqrt{\epsilon}}-\Delta(H_{S})-3\Delta(H_{S^{\prime}}),$ (S.153)
which concludes the proof. Here, we used the following conversion in the last line:
$\frac{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})|E_{S,i}-E_{S,j}|}{(\tau_{S,j}-\tau_{S,i})^{2}\frac{\epsilon+f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}{2}}$ $=\frac{2(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})\log\frac{\tau_{S,j}}{\tau_{S,i}}}{\beta(\tau_{S,j}-\tau_{S,i})^{2}(\epsilon+f(\epsilon,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j}))}$
$=\frac{2(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})\log\frac{\tau_{S,j}}{\tau_{S,i}}}{\beta(\tau_{S,j}-\tau_{S,i})^{2}(\epsilon+\frac{\epsilon\tau_{S^{\prime},i^{\prime}}+2\sqrt{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})}}{\tau_{S,i}+\tau_{S,j}-\tau_{S^{\prime},i^{\prime}}})}$
$=\frac{2(\tau_{S,i}+\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})\log\frac{\tau_{S,j}}{\tau_{S,i}}}{\beta(\tau_{S,j}-\tau_{S,i})^{2}(\epsilon(\tau_{S,i}+\tau_{S,j})+2\sqrt{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})})}$
$\geq\frac{2(\tau_{S,i}+\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})\log\frac{\tau_{S,j}}{\tau_{S,i}}}{\beta(\tau_{S,j}-\tau_{S,i})^{2}(\tau_{S,i}+\tau_{S,j}+2\sqrt{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})})}$
$=4\kappa(\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j}).$ (S.154)
VII Upper bounds for the coherence cost of general quantum operations and pairwise reversible Gibbs preserving operations (Proofs of Theorems 4 and 5)
@@
$\Lambda(\rho)=\mathrm{Tr}_{E^{\prime}}\left(V(\rho\otimes|{\eta}\rangle\!\langle{\eta}|)V^{\dagger}\right)$ (S.155)
@@
$\sqrt{{\cal F}_{c}^{\epsilon}(\Lambda)}\leq\frac{\Delta(H_{\rm tot}-V^{\dagger}H_{\rm tot}V)}{2\epsilon}+\sqrt{2}\Delta(H_{\rm tot})$ (S.156)
@@
$\sqrt{{\cal F}_{c}(\Lambda_{\epsilon})}\leq\sqrt{{\cal F}_{c}({\cal V}_{\epsilon})}\leq\frac{\Delta(H_{\rm tot}-V^{\dagger}H_{\rm tot}V)}{2\epsilon}+\sqrt{2}\Delta(H_{\rm tot}).$ (S.157)
@@
$D_{F}(\Lambda_{\epsilon},\Lambda)$ $=D_{F}(\mathrm{Tr}_{E^{\prime}}\circ{\cal V}_{\epsilon}\circ{\cal P}_{|\eta\rangle},\mathrm{Tr}_{E^{\prime}}\circ{\cal V}\circ{\cal P}_{|\eta\rangle})$ (S.158)
@@
where the second line comes from the definition of the channel purified distance (S.25), the third line is because of the data-processing inequality of the purified distance, and the fourth line follows from the assumption of ${\cal V}_{\epsilon}$. ∎
@@
$\frac{{\cal C}(\Lambda,\mathbb{P})}{\epsilon}-a\leq\sqrt{{\cal F}_{c}^{\epsilon}(\Lambda)}\leq\frac{\sqrt{2}{\cal C}(\Lambda,\mathbb{P})}{\epsilon}+a.$ (S.159)
@@
$\Lambda(\rho)=\langle{+}|{\rho}|{+}\rangle|{0}\rangle\!\langle{0}|+\langle{-}|{\rho}|{-}\rangle|{1}\rangle\!\langle{1}|.$ (S.160)
This is evidently a Gibbs-preserving Operation, noting that $H_{S^{\prime}}=0$. Take the state pair $\mathbb{P}=\{|{+}\rangle\!\langle{+}|,|{-}\rangle\!\langle{-}|\}$. Since both states are reversible under $\Lambda$, we have $\delta(\Lambda,\mathbb{P})=0$, and ${\cal C}(\Lambda,\mathbb{P})>0$ can be checked by direct computation. This ensures that Theorem S.2 can be applied, and the lower bound in (S.159) then immediately follows noting that $\Delta(H_{S})=\tilde{a}\leq a$ and $H_{S^{\prime}}=0$.
@@
$\Lambda(\rho)=\mathrm{Tr}_{S}(V\rho\otimes|{0}\rangle\!\langle{0}|_{S^{\prime}}V^{\dagger})$ (S.161)
@@
$H_{\mathrm{tot}}-V^{\dagger}H_{\mathrm{tot}}V=H_{S}\otimes\mathds{1}-V^{\dagger}(H_{S}\otimes\mathds{1})V=(H_{S}-U_{H}H_{S}U_{H})\otimes\mathds{1}.$ (S.162)
@@
$\Delta(H_{\mathrm{tot}}-V^{\dagger}H_{\mathrm{tot}}V)=\Delta(H_{S}-U_{H}H_{S}U_{H})=\sqrt{2}\tilde{a}.$ (S.163)
@@
${\cal C}(\Lambda,\mathbb{P})=\left|\langle{+}|{H_{S}-U_{H}H_{S}U_{H}}|{-}\rangle\right|=\frac{\tilde{a}}{2}.$ (S.164)
Combining (S.163) and (S.164) gives
$\Delta(H_{\mathrm{tot}}-V^{\dagger}H_{\mathrm{tot}}V)=2\sqrt{2}\,{\cal C}(\Lambda,\mathbb{P}),$ (S.165)
from which the upper bound in (S.159) follows by using Theorem S.13 and noting $\Delta(H_{S})=\tilde{a}=a/\sqrt{2}$ and $\Delta(H_{S^{\prime}})=0$. ∎